-- Main.Walth - 18 Jan 2013

We started out clear at 1" seeing, which improved to sub-arcsec for most of the night. Near the end of the A611 observation we noticed thin cirrus. At sunrise we noticed a lot more clouds than we have previously thought were there.

5:24pm: open dome

5:43pm: sunset, start spectral sky flats, counts too low (~3000 counts)! too late? 20 mins earlier?

6:00pm: finished skyflats

6:07pm: clear, wind 5m/s, clouds far far to the south

6:40pm: aligning standard, collimation looks poor

6:47pm: started integration standard, poor seeing too!

7:00pm: MACS0647 acquisition, problem with mask, appears to be stuck

7:09pm: loud bang above control room about the same time we canceled script (not sure what the noise came from)

7:39pm: MACS0647 alignment, outside temp 2.8 C, collimation issues still

7:58pm: MACS0647 start science script, seeing ~1", collimation is still bad, guide star images look triangular

9:05pm: Stopped exposure, checking collimation

9:33pm: Slew to A611 mask1

9:55pm: A611 start science script, seeing ~0.83", collimation problems still, could be mirror temperature?

11:55pm: The shell ripped (probably from 5 mins earlier)

12:01am: Resumed observing A611, seeing variable ~0.9"

12:39am: We hit the rotator limit

12:53am: Resumed A611, seeing ~0.6"

3:00am: Through slit image A611, collimation problems, probably losing flux outside of the slit

3:27am: A773 start science script, ~0.8" seeing, collimation is a significant problem, thin cirrus?

6:11am: Feige 34 start standard script, thin cirrus

6:35am: Feige 67 start standard script

7:01am: Sky was too bright for final standard
Topic revision: r5 - 18 Jan 2013, XiaohuiFan
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback